OpenStudy (anonymous):

Find the Solution of this equation -5+1/5a=6

OpenStudy (anonymous):

is the question (1/5)a or 1/5a

OpenStudy (anonymous):

@MATTW20 Here Is the question

OpenStudy (anonymous):

well we want to get a by itself. So start off by adding 5 to each side. \[-5+\frac{ 1 }{5}a+5=6+5\rightarrow \frac{ 1 }{ 5 }a=11\]

OpenStudy (anonymous):

now multiply each side by 5 \[5*\frac{ 1 }{ 5 }a=11*5\rightarrow a=55\]

OpenStudy (anonymous):

that's it btw

OpenStudy (anonymous):

Thank You for the Explanation !

OpenStudy (anonymous):

you're welcome